OpenStudy (anonymous):

Find cos θ if sin θ = -5/13 and tan θ > 0. a. -5/12 b. -13/5 c. -12/13 d. 12/5

OpenStudy (anonymous):

c

OpenStudy (zzr0ck3r):

soh cah toa o = -5 h = 13 x^2 + (-5)^2 = 13^2 x^2 = 144 so x = a = 12 cos = -12/13

OpenStudy (kenljw):

sin(x)^2 + cos(x)^2 =1 cos(x)^2 = 1- (-5/13)^2 = 144/169 cos(x)=+-12/13 because sin(x) is negative and tangent(x) is positive it must be in the third or forth quadrant. Therefore cos(x) must be negative

OpenStudy (anonymous):

c